# An Operator Splitting Method for Large-Scale CVaR-Constrained Quadratic Programs

### 07 Mar 2026 v4

@@ Abstract
We introduce a fast and scalable method for solving quadratic programs with conditional value-at-risk (CVaR) constraints. While these problems can be formulated as standard quadratic programs, the number of variables and constraints grows linearly with the number of scenarios, making general-purpose solvers impractical for large-scale problems. Our method combines operator splitting with a specialized $O(m\log m)$ algorithm for projecting onto CVaR constraints, where $m$ is the number of scenarios. The method alternates between solving a linear system and performing parallel projections: onto CVaR constraints using our specialized algorithm and onto box constraints with a closed-form solution. Numerical examples from several application domains demonstrate that our method outperforms general-purpose solvers by several orders of magnitude on problems with up to millions of scenarios. Our method is implemented in an open-source package called CVQP.
@@ 1 Introduction
Many applications in science and engineering require controlling the risk of extreme outcomes. The conditional value-at-risk (CVaR) is the expected value of losses exceeding a given quantile. Unlike variance, it penalizes only downside risk; unlike value-at-risk, it is coherent and convex. Convexity means that optimization problems involving CVaR can be solved exactly and efficiently.
@@
We present two main contributions. First, we develop an $O(m\log m)$ algorithm for projecting onto CVaR constraints, where $m$ is the number of scenarios. Building on this algorithm, our second contribution is an operator splitting method for solving large-scale CVaR-constrained quadratic programs. The method alternates between solving a linear system and performing parallel projections: one onto the CVaR constraints using our specialized algorithm and another onto box constraints with a simple closed-form solution. Numerical examples from several application domains demonstrate that our method outperforms general-purpose solvers by several orders of magnitude on problems with up to millions of scenarios.
@@ 3.3 Algorithm motivation
We want to begin decreasing elements of $v$ until the sum of the $k$ largest elements is equal to $d$. Intuitively, we should decrease the largest $k$ elements without modifying the remaining elements, since they do not affect the sum. We begin by reducing the largest elements uniformly until either the constraint is satisfied, or until the $k$th largest element ties the next element. We reduce each of the $k$ largest elements by the same amount because we are penalized by the sum of squared changes.
@@
We will need to determine this ratio at each step of the algorithm, since at a given step of the algorithm, the altered vector $v^{\prime}-\Delta$ will be partitioned into three regions: first the largest $n_{u}\in{\mbox{\bf N}}$ “untied” elements, then the $n_{t}\in{\mbox{\bf N}}$ “tied” elements, and lastly the remaining unaltered elements:
@@
The core insight of the algorithm is deriving the ratio of reducing the untied and tied elements, and is derived in §4.
@@ 4 CVaR projection algorithm
  • $a_{e}$: the value of first unaltered entry.
@@ 4.1 Decrease step
    We find the value $s_{2}\in{\mbox{\bf R}}$ such that the value of the last tied entry is equal to the value of the first unaltered entry, i.e., $z(s_{2})_{n_{u}+n_{t}}=z(s_{2})_{n_{u}+n_{t}+1}$. If $n_{u}+n_{t}=m$, there are no unaltered entries, and $s_{2}=\infty$. Otherwise, define $a_{p}$ as $\frac{n_{t}}{k-n_{u}}$ if $n_{t}=0$ and otherwise $a_{p}=1$. Define $a_{f}$ as $a_{u}$ if $n_{t}=0$ and otherwise $a_{f}=a_{t}$. Then, $s_{2}=\frac{a_{e}-a_{f}}{-a_{p}}$.
@@
    Let $s_{3}\in{\mbox{\bf R}}_{+}$ be the scaling such that the sum of the first $k$ entries of $z(s_{3})$ is equal to $d$, i.e., $s_{3}=\frac{S-d}{w}$, where $w=\sum_{i}\delta(n_{u},n_{t},k)_{i}=n_{u}(\frac{n_{t}}{k-n_{u}})+n_{t}$.
@@ 5.3 Portfolio optimization
$\begin{array}[]{ll}\text{minimize}&-\mu^{T}x+\frac{\gamma}{2}x^{T}\Sigma x\\ \text{subject to}&\mathbf{1}^{T}x=1\\ &x\geq 0\\ &\phi_{\beta}(-Rx)\leq\kappa,\end{array}$
@@ 5.4 Quantile regression
$\text{minimize}\quad\frac{1}{m}\sum_{i=1}^{m}\rho_{\tau}(y_{i}-x^{T}u_{i}-x_{0}),$
@@
$\text{minimize}\quad x^{T}\bar{u}+\phi_{\tau}(-Ux+y),$
@@
$\begin{array}[]{ll}\text{minimize}&x^{T}\bar{u}+t\\ \text{subject to}&\phi_{\tau}(-Ux+y)\leq t.\end{array}$

### 29 Mar 2026 v5

@@ Abstract
We introduce a fast and scalable method for solving quadratic programs with conditional value-at-risk (CVaR) constraints. While these problems can be formulated as standard quadratic programs, the number of variables and constraints grows linearly with the number of scenarios, making general-purpose solvers impractical for large-scale problems. Our method combines operator splitting with a specialized $O(m\log m)$ algorithm for projecting onto CVaR constraints, where $m$ is the number of scenarios. The method alternates between solving a linear system and performing parallel projections, onto CVaR constraints using our specialized algorithm and onto box constraints by simple clipping. Numerical examples from several application domains demonstrate that our method outperforms general-purpose solvers by several orders of magnitude on problems with up to millions of scenarios. Our method is implemented in an open-source package called CVQP.
@@ 1 Introduction
Many applications in finance and engineering require controlling the risk of extreme outcomes. A widely used measure for tail risk is the conditional value-at-risk (CVaR), defined as the expected value of losses exceeding a given quantile. CVaR is a coherent and convex risk measure [13], so optimization problems involving CVaR can be reliably and efficiently solved.
@@
We present two main contributions. First, we develop an $O(m\log m)$ algorithm for projecting onto CVaR constraints, where $m$ is the number of scenarios. Building on this algorithm, our second contribution is an operator splitting method for solving large-scale CVaR-constrained quadratic programs. The method alternates between solving a linear system and performing parallel projections, onto CVaR constraints using our specialized algorithm and onto box constraints by simple clipping. Numerical examples from several application domains demonstrate that our method outperforms general-purpose solvers by several orders of magnitude on problems with up to millions of scenarios.
@@ 3.3 Algorithm motivation
We want to begin decreasing elements of $v^{\prime}$ until the sum of the $k$ largest elements is equal to $d$. Intuitively, we should decrease the largest $k$ elements without modifying the remaining elements, since they do not affect the sum. We begin by reducing the largest elements uniformly until either the constraint is satisfied, or until the $k$th largest element ties the next element. We reduce each of the $k$ largest elements by the same amount because we are penalized by the sum of squared changes.
@@
We will need to determine this ratio at each step of the algorithm, since at a given step of the algorithm, the altered vector $v^{\prime}-\Delta$ will be partitioned into three regions: first the largest $n_{u}\in{\mbox{\bf N}}$ untied elements, then the $n_{t}\in{\mbox{\bf N}}$ tied elements, and lastly the remaining unaltered elements:
@@
The core insight of the algorithm is deriving the ratio of reducing the untied and tied elements; this ratio is derived in §4.
@@ 4 CVaR projection algorithm
  • $a_{e}$: the value of the first unaltered entry.
@@ 4.1 Decrease step
    We find the value $s_{2}\in{\mbox{\bf R}}$ such that the value of the last tied entry is equal to the value of the first unaltered entry, i.e., $z(s_{2})_{n_{u}+n_{t}}=z(s_{2})_{n_{u}+n_{t}+1}$. If $n_{u}+n_{t}=m$, there are no unaltered entries, and $s_{2}=\infty$. Otherwise, define $a_{p}=1$. Define $a_{f}$ as $a_{u}$ if $n_{t}=0$ and otherwise $a_{f}=a_{t}$. Then, $s_{2}=\frac{a_{e}-a_{f}}{-a_{p}}$.
@@
    Let $s_{3}\in{\mbox{\bf R}}_{+}$ be the scaling such that the sum of the first $k$ entries of $z(s_{3})$ is equal to $d$, i.e., $s_{3}=\frac{S-d}{w}$, where $w=n_{u}(\frac{n_{t}}{k-n_{u}})+(k-n_{u})$.
@@
The formulas above require $n_{t}>0$. On the first step, $n_{u}=k$ and $n_{t}=0$ (no tied entries yet), so the special-case delta applies: $\eta\leftarrow s_{0}$, $S\leftarrow S-s_{0}k$, and, if the algorithm has not terminated, $n_{t}\leftarrow 2$; the remaining updates are unchanged.
@@ 5.3 Portfolio optimization
$\begin{array}[]{ll}\mbox{minimize}&-\mu^{T}x+\frac{\gamma}{2}x^{T}\Sigma x\\ \mbox{subject to}&\mathbf{1}^{T}x=1\\ &x\geq 0\\ &\phi_{\beta}(-Rx)\leq\kappa,\end{array}$
@@ 5.4 Quantile regression
$\mbox{minimize}\quad\frac{1}{m}\sum_{i=1}^{m}\rho_{\tau}(y_{i}-x^{T}u_{i}-x_{0}),$
@@
$\mbox{minimize}\quad x^{T}\bar{u}+\phi_{\tau}(-Ux+y),$
@@
$\begin{array}[]{ll}\mbox{minimize}&x^{T}\bar{u}+t\\ \mbox{subject to}&\phi_{\tau}(-Ux+y)\leq t.\end{array}$